SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Oct. 18, 2005)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Fang Fang, Selina Li, Dong Su, Dan Walker.

Operations

  • Nothing going on for TRG at IR2 during the downtime this week.
  • There is a possibility that the DCH electronics upgrade could be delayed till next summer due to problems found in the electronics. If that was the case, we would be asked about the trigger rate in light of the high lumi we are expecting.

EMT

  • The online release expected to be done at the end of this week will contain the EMT prescale code and configuration. Rainer and Dan will work together to make a tree for this prescale parameter after the online release is installed at IR2.
  • Dan is having trouble getting the EMT turn-on curve to run in OPR. He will get help from Kris.

DCT/DCZ

  • The old DCT is back for data-taking at IR2. The essential to-do items are done:
    • Rainer updated the default crate mask to include the old DCT for all configuration types.
    • Kris installed a tag of L1TFMon V00-06-00-01 (.ps maker) and L1TMonTools V00-02-12-01 (.html for JAS) that include the old DCT for fastmon.
    • Selina put the old DCT in alarm handler for cen-mon.
  • It is confirmed that the errors such as "unphysical sl" and "bit corrupt" are results of the new TSF firmware. The new versions of the TSF firmware being tested last weekend revealed that these errors showed up in a very consistent way. The first calibration after the TSF crate power cycled was always good and produced no errors, but failed afterwards even though there was no reboot in between. Marc is working on getting another version to work. Su Dong suspected it was a problem with the DCM reset.

DCZ Online

  • Fang is working on modifying the online code so that we will ship as much corrupted data (if any) as possible to help study the corrupted data by running L1TPlayback from the xtc. Since the 3 bit tick is in, the header is packed very tightly. According to Rainer, we can ship only either input TC or module TC. Su Dong suggested to ship input TC. Rainer mentioned that a segment should at least be created so we know how much bad segments we get.

L3

  • Al wants to improve the trickle injection monitoring. He will start by improving the presentation of the existing plots under L3T page. He found out that the JAS plots were generated by a script and not in CVS. See HN posted.

Data Quality and Monitoring

  • In SP8 validation, Su Dong suspected that the L3 efficiency plots seem to indicate a problem with the denominator.
  • Su Dong also suggested to monitor the EMT prescale so if it's prescaled then we don't fill the other histograms. THe EMT->GLT transmission error histogram also needs to take this into account.
  • Rainer reminded us that the latest nightly has a T0 shift. This might be what we expected after Eric's fix, but we need his confirmation.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.